# Enhancing Topological Dependencies in Spatio-Temporal Graphs with Cycle Message Passing Blocks

### v2

@@ 7 Ablation study
We conducted an ablation study to evaluate the effect of the cycle message-passing block by testing five model variations: Cy2Mixer with the clique adjacency matrix ($A_{C}$), with the DTW matrix from PDFormer [11] for long-range dependencies, with alternative structural encodings, including Random-Walk Structural Encoding (RWSE) [35] and Laplacian Eigenvectors Encoding (LapPE) [36], and without the cycle block. Using datasets PEMS04, PEMS07, and PEMS08, results indicated that Cy2Mixer achieved the best performance when the cycle block included $A_{C}$. Additionally, generating $A_{C}$ is significantly faster than other methods like the DTW matrix, highlighting Cy2Mixer’s efficiency for spatio-temporal tasks.
@@ A.1 Dataset description
Detailed statistical information of six traffic datasets are outlined in Table 5. The first three datasets are graph-based datasets, each with a single feature, while the latter three are grid-based datasets with two features. The node counts for the grid-based datasets are 75 (15$\times$5), 270 (15$\times$18), and 1,024 (32$\times$32), respectively. The term “# Cycles" indicates the number of the cycle bases of graphs, and the term “Average Magnitude # Cycles" denotes the average number of nodes present in a cycle subgraph of each graph.
@@ A.2 Hyperparameter search
We conducted a hyperparameter search to find the optimal model, and hyperparameters for each datasets are listed in Table 6. Note that the embedding dimensions, $d_{f}$, $d_{t}$, and $d_{a}$, followed the settings of previous research [20]. The search ranges were $\{16,32,64,128\}$ for hidden dimension $d_{h}$, $\{2,3,4,5,6\}$ for number of layers, and $\{0,0.2,0.4,0.6,0.8\}$ for dropout rate, respectively. Considering previous research has demonstrated the effectiveness of incorporating tiny attention, we conducted experiments in our study to compare the performance when tiny attention is added and when it is not. The selection of the optimal model was based on its performance on the validation set. We performed experiments on STAEFormer [20] framework.
@@
In this section, a detailed comparison is conducted between the adjacency matrix and clique adjacency matrix used in Cy2Mixer, and the DTW matrix employed in PDFormer [11]. The three matrices constructed from the traffic network are visually demonstrated in Fig. 4. DTW algorithm is a similarity measure between two time-series that allows for non-linear alignment of the time series. In the context of PDFormer, DTW is used to compute the similarity of the historical traffic flow between nodes to identify the semantic neighbors of each node by calculating a pairwise distance matrix encompassing all combinations of historical traffic flow sequences. This distance matrix enables the computation of a cumulative distance matrix, representing the minimum distance between two time-series up to a specific point. Finally, the cumulative distance matrix facilitates the determination of the optimal warping path, signifying the most favorable alignment between two time-series. DTW algorithm allows the model to capture complex traffic patterns that are not easily captured by linear alignment methods, but has significant drawbacks: it is computationally time-consuming to compute and often relies on heuristic thresholds and parameters. To overcome these limitations and capture the desired spatio-temporal dependencies, we employ the use of the clique adjacency matrix. The clique adjacency matrix can be computed more efficiently than DTW since it is not dependent on time. Furthermore, it provides the model with richer topological information within the graph. The figure demonstrates that the clique adjacency matrix we use can capture topological information different from the standard adjacency matrix and DTW matrix.
Figure 4: Three matrices of PEMS04: (a) Adjacency matrix $A$, (b) Clique adjacency matrix $A_{C}$, and (c) Matrix constructed by DTW algorithm, respectively. The figure demonstrates that the clique adjacency matrix we use can capture topological information different from the standard adjacency matrix and DTW matrix.
@@ A.5 Comparison of the influence of the predefined adjacency matrix
In this section, we compared whether a predefined adjacency matrix, calculated based on the distances between nodes, makes a significant difference in the model’s performance. Recent studies in traffic prediction problems have explored not using a predefined adjacency matrix, instead learning it alongside the model [24]. To conduct this comparison, we adopted methods from previous studies to create an adjacency matrix and then generated a clique adjacency matrix from it for training in Cy2Mixer. The results of this approach are presented in Table 7. The findings indicate that Cy2Mixer does not show significant differences in performance between using a predefined $A$ and a learned $A$. Particularly, since learning $A$ and subsequently generating $A_{C}$ is costly, this study utilized a predefined $A$.
@@
The theorem is a corollary of Hatcher [39][Proposition 1.18, Corollary 2.11]. To elaborate, the map $\pi_{t_{0}}:\mathcal{G}\times I\to\mathcal{G}$ is a homotopy equivalence between the space $\mathcal{G}\times I$ representing the temporal traffic data and the space $\mathcal{G}$ representing the traffic network. The map induces an isomorphism between the fundamental groups and the homology groups of two topological spaces with rational coefficients. That is, these two maps are isomorphisms of groups for all indices $i\geq 0$:
@@
Theorem 4.1 shows in particular that any cyclic subgraph of the traffic signal $\mathcal{G}\times I$ is comprised of nodes of form $(c_{1},t_{1}),(c_{2},t_{2}),\cdots,(c_{n},t_{n})\in\mathcal{G}\times I$, where the nodes $c_{1},c_{2},\cdots,c_{n}$ are on a common cyclic subgraph of $\mathcal{G}$, and the coordinates $t_{i}$ correspond to different temporal instances. Theorem 4.1 mathematically demonstrates that the topological non-trivial invariants of a traffic network $\mathcal{G}$ can become a contributing factor for influencing the temporal variations measured among the nodes of traffic dataset, a potential aspect of temporal traffic data that may not be fully addressed from solely analyzing the set of pre-existing edges connecting the nodes of a traffic network. To elaborate, topological non-trivial invariants of $\mathcal{G}$ elucidate restrictions associated to constructing a global traffic signal $X:\mathcal{G}\times I\to\mathbb{R}^{N\times C}$ from coherently gluing a series of local temporal traffic signals $\{X_{t}\}$ measured at each time $t$. This originates from previously studied mathematical insights that obstructions in gluing continuous functions $f_{\alpha}:U_{\alpha}\to\mathbb{R}^{k}$ defined over collections of open covers $U_{\alpha}\subset Y$ of a topological space (such that $\cup_{\alpha}U_{\alpha}=Y$) to a continuous function $f:Y\to\mathbb{R}^{k}$ can be detected from topological (or cohomological) invariants of the topological space $Y$ (see for example Chapter 2 or 3 of [40]). Given that traffic forecasting beyond time $t=t_{e}$ requires a thorough understanding of traffic signals $X:\mathcal{G}\times[0,t_{e}]\to\mathbb{R}^{N\times C}$, we can hence conclude that a potential candidate to boost performances of traffic forecasting algorithms is to effectively incorporate topological invariants of traffic networks $\mathcal{G}$.
@@
As shown in Figure 6, suppose we are given with two local traffic networks (graphs) without cycles $H_{1}$ (the red graph) and $H_{2}$ (the blue graph). The dotted lines above the graphs represent an example of a traffic signal that can be measured over each local traffic network. Let $G_{1}$ and $G_{2}$ be two global traffic networks (graphs) which can be obtained from gluing the two local traffic networks in a different manner. The traffic network $G_{1}$ glues two nodes and one edge of $H_{1}$ and $H_{2}$ together (along the subnetwork notated as $A$), which does not add new cyclic subgraph to $G_{1}$. On the other hand, the traffic network $G_{2}$ glues three nodes and one edge of $H_{1}$ and $H_{2}$ together (along the subnetworks notated as $B_{1}$ and $B_{2}$), which adds a new cyclic subgraph to $G_{2}$. By gluing the two local traffic networks together, we introduce a new obstruction to traffic signals that can be measured over $G_{1}$ and $G_{2}$. The signals measured over subnetworks of local traffic network $H_{1}$ and $H_{2}$ must be identical to one another. Figure 6 clearly shows that $G_{2}$ has more obstructions on the traffic signal that can be measured in comparison to $G_{1}$. To elaborate, any traffic signal that can be measured on $G_{2}$ must satisfy the identical signals over the subnetworks $B_{1}$ and $B_{2}$, whereas $G_{1}$ only requires to satisfy the identical signals over the subnetworks $A$, which is isomorphic to the subnetwork $B_{2}$. This illustration hints that demonstrates that determining cycle structures inherent in a traffic network is crucial to understanding additional obstructions imposed over traffic signals that can be measured or predicted in the given network. Because conventional graph neural networks have limited capabilities in discerning cyclic substructures of graph datasets (as shown in [16]), adding additional architectural component which can detect cycle structures of traffic networks to the algorithm can be of crucial importance in improving accuracies in forecasting traffic signals over any traffic networks.
@@ A.7 Clique adjacency matrix
The motivation for using the clique adjacency matrix in Cy2Mixer stems from recent advances in distinguishing graph structures through the theory of covering spaces [16]. Traditional GNNs represent graphs $G$ and $H$ as identical if their universal covers—essentially infinite graphs created by unfolding the original graphs at each node—are isomorphic, and their node features match upon this unfolding. However, these universal covers inherently lack cyclic subgraphs, which are crucial structures in the original graphs. To improve the ability of GNNs to discern these cyclic features, the clique adjacency matrix ($A_{C}$) is used in Cy2Mixer. This matrix effectively transforms the geometry of the universal covers, incorporating cyclic subgraph information and thereby enhancing the model’s ability to capture and leverage topological nuances within the graph structure. The clique adjacency matrix $A_{C}$ is generated as follows: given an undirected graph $G:=(V,E)$, we first establish a cycle basis $B_{G}$ of $G$, representing a set of cyclic subgraphs that form a basis for the cycle space (or the first homology group) of $G$. The clique adjacency matrix $A_{C}$ is then constructed as the adjacency matrix of the union of complete subgraphs formed by each cycle in $B_{G}$. This is achieved by adding all possible edges among the nodes within each cycle basis element $B\in B_{G}$, enabling $A_{C}$ to represent cyclic structures comprehensively in Cy2Mixer.
@@ A.8 Performance Analysis of Cy2Mixer on the METR-LA Dataset
We also performed an experiment in the METR-LA dataset [41], which is a traffic speed dataset that records the average vehicle speeds across various sensors in Los Angeles. In our experiments on METR-LA, we evaluated the model’s performance at different forecasting horizons, specifically for time steps at horizons of 3, 6, and 12, which correspond to 15, 30, and 60 minutes into the future. The data was split into training, validation, and testing sets with ratios of 7:1:2, respectively. Table 8 presents the comparison results between our proposed method, Cy2Mixer, and various baseline models on the METR-LA dataset. Cy2Mixer demonstrates competitive performance across different forecasting horizons when compared to existing models. While Cy2Mixer does not consistently outperform all baseline models on the METR-LA dataset, it delivers results that are on par with the leading methods. The slight differences in performance metrics could be attributed to the specific characteristics of the METR-LA dataset, including its network structure and traffic patterns. Additionally, Cy2Mixer shows consistent performance across Horizons 3, 6, and 12, unlike MTGNN [24] and HIEST [27], whose performance deteriorates at Horizon 12.

### v3

@@ 7 Ablation study
We conducted an ablation study to evaluate the effect of the cycle message-passing block by testing five model variations: Cy2Mixer with the clique adjacency matrix ($A_{C}$), with the DTW matrix from PDFormer [jiang2023pdformer] for long-range dependencies, with alternative structural encodings, including Random-Walk Structural Encoding (RWSE) [dwivedi2021graph] and Laplacian Eigenvectors Encoding (LapPE) [rampavsek2022recipe], and without the cycle block. Using datasets PEMS04, PEMS07, and PEMS08, results indicated that Cy2Mixer achieved the best performance when the cycle block included $A_{C}$. Additionally, generating $A_{C}$ is significantly faster than other methods like the DTW matrix, highlighting Cy2Mixer’s efficiency for spatio-temporal tasks.
@@ A.1 Dataset description
[…]
@@ A.2 Hyperparameter search
We conducted a hyperparameter search to find the optimal model, and hyperparameters for each datasets are listed in Table 6. Note that the embedding dimensions, $d_{f}$, $d_{t}$, and $d_{a}$, followed the settings of previous research [liu2023spatio]. The search ranges were $\{16,32,64,128\}$ for hidden dimension $d_{h}$, $\{2,3,4,5,6\}$ for number of layers, and $\{0,0.2,0.4,0.6,0.8\}$ for dropout rate, respectively. Considering previous research has demonstrated the effectiveness of incorporating tiny attention, we conducted experiments in our study to compare the performance when tiny attention is added and when it is not. The selection of the optimal model was based on its performance on the validation set. We performed experiments on STAEFormer [liu2023spatio] framework.
@@
In this section, a detailed comparison is conducted between the adjacency matrix and clique adjacency matrix used in Cy2Mixer, and the DTW matrix employed in PDFormer [jiang2023pdformer]. The three matrices constructed from the traffic network are visually demonstrated in Fig. 4. DTW algorithm is a similarity measure between two time-series that allows for non-linear alignment of the time series. In the context of PDFormer, DTW is used to compute the similarity of the historical traffic flow between nodes to identify the semantic neighbors of each node by calculating a pairwise distance matrix encompassing all combinations of historical traffic flow sequences. This distance matrix enables the computation of a cumulative distance matrix, representing the minimum distance between two time-series up to a specific point. Finally, the cumulative distance matrix facilitates the determination of the optimal warping path, signifying the most favorable alignment between two time-series. DTW algorithm allows the model to capture complex traffic patterns that are not easily captured by linear alignment methods, but has significant drawbacks: it is computationally time-consuming to compute and often relies on heuristic thresholds and parameters. To overcome these limitations and capture the desired spatio-temporal dependencies, we employ the use of the clique adjacency matrix. The clique adjacency matrix can be computed more efficiently than DTW since it is not dependent on time. Furthermore, it provides the model with richer topological information within the graph. The figure demonstrates that the clique adjacency matrix we use can capture topological information different from the standard adjacency matrix and DTW matrix.
Figure 4: Three matrices of PEMS04: (a) Adjacency matrix AA, (b) Clique adjacency matrix $A_{C}$, and (c) Matrix constructed by DTW algorithm, respectively. The figure demonstrates that the clique adjacency matrix we use can capture topological information different from the standard adjacency matrix and DTW matrix.
@@ A.5 Comparison of the influence of the predefined adjacency matrix
In this section, we compared whether a predefined adjacency matrix, calculated based on the distances between nodes, makes a significant difference in the model’s performance. Recent studies in traffic prediction problems have explored not using a predefined adjacency matrix, instead learning it alongside the model [wu2020connecting]. To conduct this comparison, we adopted methods from previous studies to create an adjacency matrix and then generated a clique adjacency matrix from it for training in Cy2Mixer. The results of this approach are presented in Table 7. The findings indicate that Cy2Mixer does not show significant differences in performance between using a predefined $A$ and a learned $A$. Particularly, since learning $A$ and subsequently generating $A_{C}$ is costly, this study utilized a predefined $A$.
@@
The theorem is a corollary of Ha02[Proposition 1.18, Corollary 2.11]. To elaborate, the map $\pi_{t_{0}}:\mathcal{G}\times I\to\mathcal{G}$ is a homotopy equivalence between the space $\mathcal{G}\times I$ representing the temporal traffic data and the space $\mathcal{G}$ representing the traffic network. The map induces an isomorphism between the fundamental groups and the homology groups of two topological spaces with rational coefficients. That is, these two maps are isomorphisms of groups for all indices $i\geq 0$:
@@
Theorem 4.1 shows in particular that any cyclic subgraph of the traffic signal $\mathcal{G}\times I$ is comprised of nodes of form $(c_{1},t_{1}),(c_{2},t_{2}),\cdots,(c_{n},t_{n})\in\mathcal{G}\times I$, where the nodes $c_{1},c_{2},\cdots,c_{n}$ are on a common cyclic subgraph of $\mathcal{G}$, and the coordinates $t_{i}$ correspond to different temporal instances. Theorem 4.1 mathematically demonstrates that the topological non-trivial invariants of a traffic network $\mathcal{G}$ can become a contributing factor for influencing the temporal variations measured among the nodes of traffic dataset, a potential aspect of temporal traffic data that may not be fully addressed from solely analyzing the set of pre-existing edges connecting the nodes of a traffic network. To elaborate, topological non-trivial invariants of $\mathcal{G}$ elucidate restrictions associated to constructing a global traffic signal $X:\mathcal{G}\times I\to\mathbb{R}^{N\times C}$ from coherently gluing a series of local temporal traffic signals $\{X_{t}\}$ measured at each time $t$. This originates from previously studied mathematical insights that obstructions in gluing continuous functions $f_{\alpha}:U_{\alpha}\to\mathbb{R}^{k}$ defined over collections of open covers $U_{\alpha}\subset Y$ of a topological space (such that $\cup_{\alpha}U_{\alpha}=Y$) to a continuous function $f:Y\to\mathbb{R}^{k}$ can be detected from topological (or cohomological) invariants of the topological space $Y$ (see for example Chapter 2 or 3 of [Ha77]). Given that traffic forecasting beyond time $t=t_{e}$ requires a thorough understanding of traffic signals $X:\mathcal{G}\times[0,t_{e}]\to\mathbb{R}^{N\times C}$, we can hence conclude that a potential candidate to boost performances of traffic forecasting algorithms is to effectively incorporate topological invariants of traffic networks $\mathcal{G}$.
@@
As shown in Figure 6, suppose we are given with two local traffic networks (graphs) without cycles $H_{1}$ (the red graph) and $H_{2}$ (the blue graph). The dotted lines above the graphs represent an example of a traffic signal that can be measured over each local traffic network. Let $G_{1}$ and $G_{2}$ be two global traffic networks (graphs) which can be obtained from gluing the two local traffic networks in a different manner. The traffic network $G_{1}$ glues two nodes and one edge of $H_{1}$ and $H_{2}$ together (along the subnetwork notated as $A$), which does not add new cyclic subgraph to $G_{1}$. On the other hand, the traffic network $G_{2}$ glues three nodes and one edge of $H_{1}$ and $H_{2}$ together (along the subnetworks notated as $B_{1}$ and $B_{2}$), which adds a new cyclic subgraph to $G_{2}$. By gluing the two local traffic networks together, we introduce a new obstruction to traffic signals that can be measured over $G_{1}$ and $G_{2}$. The signals measured over subnetworks of local traffic network $H_{1}$ and $H_{2}$ must be identical to one another. Figure 6 clearly shows that $G_{2}$ has more obstructions on the traffic signal that can be measured in comparison to $G_{1}$. To elaborate, any traffic signal that can be measured on $G_{2}$ must satisfy the identical signals over the subnetworks $B_{1}$ and $B_{2}$, whereas $G_{1}$ only requires to satisfy the identical signals over the subnetworks $A$, which is isomorphic to the subnetwork $B_{2}$. This illustration hints that demonstrates that determining cycle structures inherent in a traffic network is crucial to understanding additional obstructions imposed over traffic signals that can be measured or predicted in the given network. Because conventional graph neural networks have limited capabilities in discerning cyclic substructures of graph datasets (as shown in [choi2022cycle]), adding additional architectural component which can detect cycle structures of traffic networks to the algorithm can be of crucial importance in improving accuracies in forecasting traffic signals over any traffic networks.
@@ A.7 Clique adjacency matrix
The motivation for using the clique adjacency matrix in Cy2Mixer stems from recent advances in distinguishing graph structures through the theory of covering spaces [choi2022cycle]. Traditional GNNs represent graphs $G$ and $H$ as identical if their universal covers—essentially infinite graphs created by unfolding the original graphs at each node—are isomorphic, and their node features match upon this unfolding. However, these universal covers inherently lack cyclic subgraphs, which are crucial structures in the original graphs. To improve the ability of GNNs to discern these cyclic features, the clique adjacency matrix ($A_{C}$) is used in Cy2Mixer. This matrix effectively transforms the geometry of the universal covers, incorporating cyclic subgraph information and thereby enhancing the model’s ability to capture and leverage topological nuances within the graph structure. The clique adjacency matrix $A_{C}$ is generated as follows: given an undirected graph $G:=(V,E)$, we first establish a cycle basis $B_{G}$ of $G$, representing a set of cyclic subgraphs that form a basis for the cycle space (or the first homology group) of $G$. The clique adjacency matrix $A_{C}$ is then constructed as the adjacency matrix of the union of complete subgraphs formed by each cycle in $B_{G}$. This is achieved by adding all possible edges among the nodes within each cycle basis element $B\in B_{G}$, enabling $A_{C}$ to represent cyclic structures comprehensively in Cy2Mixer.
@@ A.8 Performance Analysis of Cy2Mixer on the METR-LA Dataset
We also performed an experiment in the METR-LA dataset [chen2001freeway], which is a traffic speed dataset that records the average vehicle speeds across various sensors in Los Angeles. In our experiments on METR-LA, we evaluated the model’s performance at different forecasting horizons, specifically for time steps at horizons of 3, 6, and 12, which correspond to 15, 30, and 60 minutes into the future. The data was split into training, validation, and testing sets with ratios of 7:1:2, respectively. Table 8 presents the comparison results between our proposed method, Cy2Mixer, and various baseline models on the METR-LA dataset. Cy2Mixer demonstrates competitive performance across different forecasting horizons when compared to existing models. While Cy2Mixer does not consistently outperform all baseline models on the METR-LA dataset, it delivers results that are on par with the leading methods. The slight differences in performance metrics could be attributed to the specific characteristics of the METR-LA dataset, including its network structure and traffic patterns. Additionally, Cy2Mixer shows consistent performance across Horizons 3, 6, and 12, unlike MTGNN [wu2020connecting] and HIEST [ma2023rethinking], whose performance deteriorates at Horizon 12.